### v2

# GIFT: Guided Importance-Aware Fine-Tuning for Diffusion Language Models

Abstract
Diffusion models have recently shown strong potential in language modeling, offering faster generation compared to traditional autoregressive approaches. However, applying supervised fine-tuning (SFT) to diffusion models remains challenging, as they lack precise probability estimates at each denoising step. While the diffusion mechanism enables the model to reason over entire sequences, it also makes the generation process less predictable and often inconsistent. This highlights the importance of controlling key tokens that guide the direction of generation. To address this issue, we propose GIFT, an importance-aware finetuning method for diffusion language models, where tokens are assigned different importance weights based on their entropy. Derived from diffusion theory, GIFT delivers substantial gains: across diverse settings including different mainstream training datasets ranging from 1k to 10k in size, utilizing LoRA or full parameter fine-tuning, and training on base or instruct models, GIFT consistently achieves superior overall performance compared to standard SFT on four widely used reasoning benchmarks (Sudoku, Countdown, GSM8K, and MATH-500).
GIFT: Guided Importance-Aware Fine-Tuning for Diffusion Language Models
Guowei Xu†, Wenxin Xu†, Jiawang Zhao, Kaisheng Ma* Institute for Interdisciplinary Information Sciences Tsinghua University †Equal Contribution. Corresponding Author: kaisheng@tsinghua.edu.cn
1 Introduction
Recently, the landscape of large language modeling has been reshaped by the emergence of diffusion-based approaches  (Nie et al., 2025; Ou et al., 2025; Yang et al., 2025), which offer a fundamentally different perspective on sequence generation compared to autoregressive (AR) methods  Radford et al. (2018, 2019); Brown et al. (2020). Rather than relying on strictly sequential decoding, diffusion large language models (dLLMs) employ iterative refinement procedures that enable efficient parallel generation and open new possibilities for scaling language models.
Currently, similar to AR models, the training paradigm of dLLMs is generally divided into three stages: pre-training, instruction tuning, and reinforcement learning (RL). Among them, the pre-training and instruction tuning stages adopt an SFT loss specifically derived for dLLMs (Nie et al., 2025), while the RL stage employs a loss estimated on the basis of the PPO algorithm (Schulman et al., 2017) and tailored to the characteristics of dLLMs  (Zhu et al., 2025).
Specifically, when designing the SFT objective, Nie et al., 2025 model the generation process of dLLMs as a continuous-time diffusion process, where the diffusion timestep $t$ ranges within $[0,1]$. Here, $t=0$ denotes a completely original, non-masked sequence, and $t=1$ denotes a fully masked sequence. As illustrated in Figure 1(a), let the original response of the model be $x_{0}$. At a given timestep $t$, each token is masked with probability $t$, resulting in a partially masked response $x_{t}$. The training objective of dLLMs is then to enable the model to reconstruct the masked parts of $x_{0}$ given the partially masked input $x_{t}$. Based on this formulation, the loss is defined as follows:
$L=-\sum_{i}\mathbb{E}_{t}\left[\mathbf{1}[x_{t}^{i}=\mathbf{M}]\,\frac{1}{t}\log p\big(x_{0}^{i}\mid x_{t}\big)\right]$ (1)
This design of the SFT loss is intuitive and aligns well with the underlying rationale of diffusion theory. However, it presents several critical limitations. Among them, one of the most prominent issues is that the loss implicitly assumes a uniform masking rate across all tokens, thereby treating each token as equally important throughout the diffusion process. This assumption overlooks the inherent heterogeneity in token significance: tokens that play a central role in planning and reasoning should arguably receive greater emphasis during training than other tokens.
To address this discrepancy, we introduce GIFT, a guided importance-aware finetuning mechanism based on predictive uncertainty. Specifically, we find that the entropy of the model’s predictive distribution over tokens serves as a reliable metric for token importance at the current training stage. Tokens associated with higher entropy correspond to positions where the model exhibits greater uncertainty in generation. By prioritizing these high-entropy tokens during training via an adaptive importance weighting strategy applied to the SFT loss, we can concentrate computation and importance on tokens where the model is currently more uncertain, thereby improving the overall training effectiveness.
Figure 1: (a) The SFT pipeline. A timestep $t$ is uniformly sampled from $[0,1]$, and each token is masked independently with probability $t$. The training objective is to predict the masked tokens accurately based on the unmasked ones. (b) The GIFT pipeline. In each training step, we perform two forward passes. During the first forward pass, we mask the entire answer and estimate the masking rate $\beta_{i}$ for each token by computing its predictive entropy. In the second forward pass, the $i$-th token is masked with probability $t_{i}$ (computed from $\beta_{i}$), and its training weight is set to $\tfrac{1}{t_{i}}$. Tokens with higher entropy are more likely to be masked and thus receive stronger training signals.
Based on the fundamental principles of diffusion models, we derive an importance-weighted SFT loss formulation that is theoretically consistent with diffusion dynamics. Empirically, we evaluate this approach on four widely used benchmarks, Sudoku (Arel, 2025), Countdown (Pan et al., 2025), GSM8K (Cobbe et al., 2021), and MATH500 (Lightman et al., 2024), by comparing models fine-tuned with the standard SFT algorithm against those fine-tuned with GIFT under identical training data. Specifically, our method is trained using LoRA (Hu et al., 2022) on s1K, s1K-1.1 (Muennighoff et al., 2025), and 3k samples drawn from the Mixture-of-Thoughts dataset of open-r1 (HuggingFace, 2025), and using full parameter fine-tuning on 10k samples drawn from Tulu3 (Lambert et al., 2025). Across these diverse training settings, our approach consistently achieves superior overall performance compared to standard SFT.
In addition, ablation studies further confirm that both the theoretically derived importance-weighted SFT loss and the entropy-based importance-weighting scheme are indispensable for achieving the observed performance improvements. Finally, the time efficiency analysis shows that GIFT incurs only minimal overhead during LoRA fine-tuning, and imposes no computational overhead during full-parameter fine-tuning, even achieving faster speeds than standard SFT.
In summary, our core contributions are as follows:
  • We derive a theoretically grounded formulation of the importance-weighted SFT loss from the perspective of diffusion processes, ensuring consistency with the underlying principles of diffusion-based generation.
  • We identify token-level predictive entropy as a reasonable and effective metric for capturing token importance, and we demonstrate both theoretically and empirically that tokens with higher entropy should be emphasized during training to enhance the effectiveness and stability of the training process.
  • Building upon these insights, we propose GIFT, an importance-weighted finetuning algorithm that consistently achieves superior overall performance compared to the standard SFT approach across multiple training sets and benchmarks. Ablation studies further validate the necessity and effectiveness of each design choice in GIFT. Time efficiency analysis demonstrates that GIFT incurs minimal training time overhead compared to SFT.
2 Preliminaries: Continuous-Time Discrete Diffusion Model
Following  Ou et al., 2025, we model diffusion language models as a Continuous-Time Discrete Diffusion Model (Sun et al., 2023). Intuitively, in each infinitesimal time interval $\Delta t$, every token has a certain probability of being masked.
Formally, this process can be described as a time-dependent Markov chain, whose transition dynamics are governed by a matrix $Q_{t}$. For a transition from state $x$ to $y$ within $(t,t+\Delta t)$, we have
$p_{t+\Delta t|t}(y|x)=\left\{\begin{aligned} &Q_{t}(x,y)\Delta t+o(\Delta t),\\ &\quad\text{if }x\neq y\\ &1+Q_{t}(x,x)\Delta t+o(\Delta t),\\ &\quad\text{if }x=y\end{aligned}\right$. (2)
where $Q_{t}(x,y)\geq 0$ for $x\neq y$ and $Q_{t}(x,x)<0$. Thus, $Q_{t}(x,y)$ specifies the instantaneous transition rate from state $x$ to $y$.
In practice, it is common to assume a time–factorized form (Campbell et al., 2022), namely
$Q_{t}=f(t)Q,$ (3)
where $Q$ is a constant matrix and $f(t)$ is a scalar function controlling the evolution speed.
Let $P_{t|s}(x,y)=p_{t|s}(y|x)$ denote the transition probability matrix from time $s$ to $t$. Then, $P_{t|s}$ satisfies the Kolmogorov’s forward equation (Anderson, 2012):
$\frac{d}{dt}P_{t|s}=f(t)P_{t|s}Q.$ (4)
The solution is given by
$P_{t|s}=\exp\!\left(Q\int_{s}^{t}f(u)\,du\right),$ (5)
where $\exp$ denotes the matrix exponential.
For notational convenience, we define
$\bar{f}(x)=\int_{0}^{x}f(t)\,dt.$ (6)
With this shorthand, the solution can be equivalently written as
$P_{t|s}=\exp\!\big((\bar{f}(t)-\bar{f}(s))Q\big).$ (7)
Intuitively, $s$ corresponds to the sequence before masking at the current step, $t$ corresponds to the sequence after masking at the current step, and $Q$ represents the masking rate of each token. In prior work, it is commonly assumed that the $Q$ matrix takes the following form:  (Ou et al., 2025; Nie et al., 2025)
$Q=\begin{bmatrix}-1&0&\cdots&0&1\\ 0&-1&\cdots&0&1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-1&1\\ 0&0&\cdots&0&0\end{bmatrix}$ (8)
Under this assumption, all tokens gradually diffuse into an absorbing state [M], that is, they become masked.
Frequent tokens with high entropy
Frequent tokens with low entropy
Figure 2: Visualization of high-frequency tokens with different entropy levels. Tokens with higher entropy correspond to those where the model exhibits greater uncertainty, and are learned more frequently during training.
3 Method
3.1 Overview
We propose GIFT (Guided Importance-Aware Fine-Tuning), an importance-weighted finetuning method designed to assign higher importance weights to tokens where the current model exhibits greater uncertainty. As we will explain later, in practice, we quantify token importance using the square root of token entropy.
As shown in Figure 1(a), the original SFT approach masks all tokens with the same probability $t$, and each token contributes equally to the loss function with a weight of $\frac{1}{t}$. In contrast, Figure 1(b) illustrates the GIFT pipeline. In each training step, we perform two forward passes. In the first pass, the entire answer is masked, and the masking rate $\beta_{i}$ for each token is estimated based on its predictive entropy. In the second pass, each token is independently masked with probability $t_{i}$ (computed from $\beta_{i}$), and its training weight is assigned as $\tfrac{1}{t_{i}}$. This design ensures that tokens with higher entropy are more likely to be masked and trained more frequently, thereby improving the effectiveness and stability of the training process.
The derivation of the loss function is presented in Section 3.2, while the details and motivation of the importance weighting scheme are provided in Section 3.3 and Appendix D. A comprehensive description of the implementation of GIFT can be found in Appendix B.
3.2 Importance-aware Loss Derivation
As noted earlier, prior work typically assumes that all tokens are masked at the same rate, leading to a $Q$ matrix that takes the form of an absorbing matrix composed of $1$ and 1-1. In contrast, we assign each token a distinct masking rate $\beta$, from which we derive an importance-weighted finetuning formulation. Importantly, the resulting SFT loss remains consistent with the fundamental properties of diffusion, since the $Q$ matrix is still well-defined under this construction.
Namely, we now set $Q$ as:
$Q=\begin{bmatrix}-\beta_{1}&0&\cdots&0&\beta_{1}\\ 0&-\beta_{2}&\cdots&0&\beta_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-\beta_{n-1}&\beta_{n-1}\\ 0&0&\cdots&0&0\end{bmatrix}$ (9)
Based on the newly defined form of the $Q$ matrix, and following the derivation in  Ou et al., 2025, we obtain the modified formulation of the SFT loss. The detailed derivation is provided in Appendix A.
Theorem 1.
Assuming the $Q$ matrix takes the form given in Equation 9, let the initial sequence be $x_{0}$ and the sequence at time $t$ be $x_{t}$. Under this setting, the $i$-th token is masked with probability $t_{i}=1-(1-t)^{\tfrac{\beta_{x^{i}}}{\beta_{\text{ref}}}}$, where $\beta_{x^{i}}$ denotes the masking rate of the $i$-th token, and $\beta_{\text{ref}}$ is a specified reference masking rate. Moreover, the importance-weighted supervised finetuning loss can be derived as follows:
$L=-\sum_{i}\mathbb{E}_{t_{i}}\left[\mathbf{1}\!\left[x_{t_{i}}^{i}=\mathbf{M}\right]\frac{1}{t_{i}}\log p\big(x_{0}^{i}\mid x_{t}\big)\right]$ (10)
Intuitively, under the importance-weighted supervised finetuning loss, a token $x_{i}$ with a larger masking rate $\beta_{x^{i}}$ corresponds to a larger $t_{i}$, making it more likely to be masked and subsequently learned. Here, $\beta_{\mathrm{ref}}$ can be any reference value. Empirically, we set $\beta_{\mathrm{ref}}$ to the mean of all $\beta$ values to enhance numerical stability.
3.3 Motivation of Using Entropy as the Masking Rate Metric
Having established the form of the importance-weighted supervised finetuning loss, we next investigate how to assign masking rate to tokens. First, we require that the chosen metric does not introduce excessive computational overhead. In particular, it is infeasible to compute such a metric for every token individually, and we can only afford $O(1)$ additional forward passes in total. For this reason, we consider masking all the answers and then examining the distribution of the model’s output logits $z_{i}$ to be a practical choice. Specifically, given a problem $x$ and its answer $y$, we mask the entire answer and compute
$z_{i}=\text{model}(\cdot\mid x,[M]).\text{logits}.$ (11)
We then determine the masking rate $\beta_{i}$ of the $i$-th token. Based on our analysis, the entropy of the distribution of $z_{i}$ serves as an effective metric. First, in Appendix D, drawing inspiration from theoretical insights, we demonstrate that using entropy as an importance weight may help reduce the variance of the loss estimation. Furthermore, as shown in Figure 2, we empirically visualize the entropy distribution among the 100 most frequent tokens during training, contrasting tokens with higher versus lower entropy. It can be observed that tokens with higher entropy tend to carry richer information and play a more critical role in the model’s decision-making. For example, words such as “first” and “second” exhibit higher entropy, often signaling logical or sequential structures, whereas words like “all” and “such” show lower entropy and function primarily as auxiliary words, contributing less to semantic content.
Since we empirically find that using the raw entropy values may destabilize training (Section  4.3.3), we use the square root of entropy, which yields more stable optimization in practice:
$\beta_{i}=\sqrt{\,H(\mathrm{softmax}(z_{i}))\,}.$ (12)
4 Experiment
In this section, we first describe the experimental setup, followed by a comparison of the performance of GIFT and the standard SFT algorithm across multiple training datasets and benchmarks. We then conduct ablation studies to validate the necessity of each design component in GIFT, and finally provide a runtime comparison between GIFT and SFT under identical settings.
Model Sudoku Countdown GSM8K MATH500 Average
0-shot 0-shot 0-shot 0-shot
Base Instruct Model
LLaDA-8B-Instruct 5.5 16.0 76.7 32.4 32.6
s1K Dataset, LoRA Tuning
Instruct Model + SFT 4.6 (-0.9) 23.8 (+7.8) 78.8 (+2.1) 32.6 (+0.2) 34.9 (+2.3)
Instruct Model + GIFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
s1K-1.1 Dataset, LoRA Tuning
Instruct Model + SFT 4.2 (-1.3) 21.5 (+5.5) 78.4 (+1.7) 30.8 (-1.6) 33.7 (+1.1)
Instruct Model + GIFT 6.8 (+1.3) 20.3 (+4.3) 78.1 (+1.4) 32.6 (+0.2) 34.4 (+1.8)
openr1-3k Dataset, LoRA Tuning
Instruct Model + SFT 6.3 (+0.8) 17.2 (+1.2) 77.3 (+0.6) 32.0 (-0.4) 33.2 (+0.6)
Instruct Model + GIFT 6.9 (+1.4) 17.6 (+1.6) 77.5 (+0.8) 32.8 (+0.4) 33.7 (+1.1)
Tulu3-10k Dataset, Full Paramter Tuning
Base Model + SFT 5.0 18.7 66.1 23.2 28.2
Base Model + GIFT 4.8 20.3 65.4 25.0 28.9
Table 1: Comparison of the performance of GIFT and SFT across different training datasets and benchmarks. The table shows that GIFT achieves substantial percentage improvements over SFT.
4.1 Experimental Setup
We adopt LLaDA-8B-Base and LLaDA-8B-Instruct (Nie et al., 2025) as base models and apply either LoRA fine-tuning (Hu et al., 2022) or full parameter fine-tuning. For the training datasets, we use s1K, s1K-1.1, openr1-3k, and Tulu3-10k to evaluate the performance of GIFT and SFT. Among them, s1K and s1K-1.1 are reasoning datasets proposed by Muennighoff et al., 2025, while openr1-3k is constructed by sampling 3k examples from the Mixture-of-Thoughts dataset of open-r1 (HuggingFace, 2025), with sequence lengths less than 2.5k under the open-r1 encoder. Tulu3-10k consists of 10k samples drawn from Tulu3 (Lambert et al., 2025), following the setting of Zhou et al., 2025. To verify that our algorithm is framework-agnostic, we employ the wd1  (Tang et al., 2025) framework for LoRA training, and the dllm (Zhou et al., 2025) framework for full-parameter fine-tuning.
The evaluation is conducted on the following benchmarks:
  • Sudoku (Arel, 2025): A symbolic reasoning benchmark requiring models to solve Sudoku puzzles through step-by-step logical deduction.
  • Countdown (Pan et al., 2025): A numerical reasoning dataset where models must combine given numbers using arithmetic operations to reach a target value.
  • GSM8K (Cobbe et al., 2021): A widely used benchmark of grade-school math word problems designed to evaluate arithmetic and reasoning abilities.
  • MATH500 (Lightman et al., 2024): A subset of the MATH dataset containing 500 challenging competition-level math problems that test advanced mathematical reasoning.
The detailed training and evaluation hyperparameters and settings are provided in Appendix C.
4.2 Main Results
As shown in Table 1, we compare the performance of GIFT and the standard SFT method across different training datasets and benchmarks. Due to computational constraints, each of our experiments is the result of a single run. Notably, GIFT consistently achieves higher average accuracy than SFT, demonstrating that our method achieves effective training by leveraging token entropy. Experiments across datasets of varying sizes, different training frameworks, multiple benchmarks, and on both base and instruct models demonstrate that the advantage of GIFT over SFT is universal rather than case-specific, indicating the strong generalizability of our method.
4.3 Ablation Study
In this subsection, we conduct ablation studies to validate the design choices of GIFT, focusing on three aspects: the necessity of using entropy as the metric, the validity of the theoretically derived loss function, and the empirical necessity of applying the square root of entropy during training.
4.3.1 Necessity of Using Entropy as the Metric
Model Sudoku Countdown GSM8K MATH500 Average
0-shot 0-shot 0-shot 0-shot
Base Model
LLaDA-8B-Instruct 5.5 16.0 76.7 32.4 32.6
s1K Dataset
+ SFT 4.6 (-0.9) 23.8 (+7.8) 78.8 (+2.1) 32.6 (+0.2) 34.9 (+2.3)
+ Weighted SFT (- log p) 3.0 (-2.5) 16.4 (+0.4) 75.4 (-1.3) 29.2 (-3.2) 31.0 (-1.6)
+ GIFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
Table 2: Comparison between entropy-based importance weighting and NLL-based importance weighting. As shown in the table, the SFT algorithm with NLL-based importance weighting performs significantly worse than the entropy-based variant, and on many tasks even underperforms the original model.
To demonstrate the necessity of entropy-based importance weighting, we compared our method with a baseline using the negative log-likelihood (NLL). Unlike entropy-based importance weighting, NLL emphasizes tokens that appear less frequently rather than those that are difficult to predict or carry more information. Assuming the logits are $z_{i}$, the masking rate $\beta_{i}$ for the $i$-th token under the NLL setting is:
$\beta_{i}=-\log\bigl(\text{softmax}(z_{i})^{x_{0}^{i}}\bigr)$ (13)
As shown in Table 2, the SFT method weighted by NLL performs poorly, in some cases even worse than the original base model. This underscores the importance of focusing on high-entropy, informative tokens for achieving effective fine-tuning performance.
4.3.2 Validity of the Theoretically Derived Loss Function
Our loss function is derived from the continuous-time discrete diffusion model combined with the Denoising Score Entropy (DSE) loss. As a result, our model preserves the inherent characteristics of the diffusion process more effectively than methods that simply add a weight to the loss function:
$L=-\sum_{i}\mathbb{E}_{t}\left[\mathbf{1}\!\left[x_{t}^{i}=\mathbf{M}\right]\frac{w_{i}}{t}\log p\left(x_{0}^{i}\mid x_{t}\right)\right]$ (14)
To illustrate this, we compare our approach with two weighted SFT methods by applying a simple token-wise weighting, namely using our square-root entropy as the weight, and using the Dream (Ye et al., 2025) method. Specifically, when using our square-root entropy, the weight of the $i$-th token is:
$w_{i}=\sqrt{\,H(\mathrm{softmax}(z_{i}))\,}$ (15)
When using the weighting method from Dream, the weight of the $i$-th token is:
$w_{i}=\tfrac{1}{2}\sum_{j=1}^{N}\mathbf{1}\!\left[x_{t}^{j}\neq\mathbf{M}\right]\,\mathrm{Geo}\!\left(p,\lvert j-i\rvert-1\right)$ (16)
Here, Geo denotes the geometric distribution, and $p$ controls the sharpness of the distribution. Since the original work did not provide the corresponding parameter setting, we set it to a moderate value of 0.3.
Model Sudoku Countdown GSM8K MATH500 Average
0-shot 0-shot 0-shot 0-shot
Base Model
LLaDA-8B-Instruct 5.5 16.0 76.7 32.4 32.6
s1K Dataset
+ SFT 4.6 (-0.9) 23.8 (+7.8) 78.8 (+2.1) 32.6 (+0.2) 34.9 (+2.3)
+ Dream 3.3 (-2.2) 21.1 (+5.1) 77.0 (+0.3) 30.0 (-2.4) 32.8 (+0.2)
+ GIFT (SW) 5.3 (-0.2) 23.8 (+7.8) 75.9 (-0.8) 31.6 (-0.8) 34.2 (+1.6)
+ GIFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
Table 3: Comparison between GIFT, GIFT (SW) (simple token-wise weighting using our square-root entropy as the weight), and Dream. Dream yields negligible improvements. GIFT (SW) provides some gains, but still performs worse than SFT, whereas GIFT significantly outperforms SFT.
Table 3 presents a comparison between GIFT, GIFT (SW) (i.e., simple token-wise weighting using our square-root entropy as the weight), and Dream. Dream yields only minimal improvements and, in some cases, performs worse than standard SFT. GIFT (SW) provides some gains but still performs worse than SFT, whereas GIFT significantly outperforms SFT. This demonstrates that the diffusion loss derived from theoretical principles possesses superior properties.
4.3.3 Empirical Necessity of Applying the Square Root of Entropy During Training
Below, we explain why we choose the square root of entropy rather than entropy as the metric. As shown in Table 4, we observed that directly using token weights based on raw entropy leads to large variations across tokens, which can significantly reduce training stability and even cause oscillations. To address this issue, we implement GIFT using the square root of entropy as the weighting factor. This modification effectively reduces the scale of the gradient norms, resulting in more stable training and improved overall convergence behavior.
Gradient Norm Entropy $\sqrt{\text{Entropy}}$
Max 646.1 46.2
Median 0.47 0.43
Average 2.02 1.03
Table 4: Comparison of gradient norms during training when using entropy versus square root entropy. Using entropy sometimes leads to extremely large gradient norms, resulting in instability, whereas using square root entropy effectively alleviates this issue.
4.4 Time Efficiency Analysis
In the time efficiency analysis, there are two counteracting factors. First, GIFT requires an additional forward pass compared to SFT, resulting in a slightly higher computational overhead. However, it is important to note that the cost of a single gradient-free forward pass is significantly lower than that of a backward pass; thus, this factor only causes a marginal decrease in speed. Second, the expected number of masked tokens per iteration may differ between GIFT and SFT, which also impacts the overall training speed.
Dataset Framework SFT GIFT
s1K wd1 46 min 57 min
Tulu3 dllm 28 min 26 min
Table 5: Time efficiency comparison between SFT and GIFT on different tasks and frameworks.
As shown in Table 5, we compare the training speed of GIFT and SFT under two distinct settings: (1) LoRA fine-tuning on the s1K dataset for 20 epochs using the wd1 training framework on 4 H100 GPUs; and (2) full fine-tuning on the Tulu3-10k dataset for 5 epochs using the dllm training framework on 8 H100 GPUs. These scenarios represent representative results selected from our main experiments. The results indicate that while GIFT is merely slightly slower than SFT during LoRA fine-tuning, it is marginally faster than SFT during full fine-tuning. We attribute this to the fact that the additional forward pass incurs only a minor overhead. Conversely, since GIFT prioritizes more significant tokens, the expected total number of masked tokens may be lower than that of SFT, thereby reducing the computational load during backpropagation. This reduction in computation has a more pronounced impact during full fine-tuning compared to LoRA, leading to the observation that our method achieves faster speeds than SFT in the full fine-tuning setting.
5 Related Work
5.1 Diffusion Large Language Models
Diffusion models, initially developed for continuous data, have recently been adapted to discrete sequences, including natural language. Early work such as Structured Denoising Diffusion Models (D3PMs) (Austin et al., 2021) extended discrete diffusion to text, showing competitive performance in character-level generation and large vocabulary language modeling. DiffusionBERT (He et al., 2023) combined masked language modeling with diffusion, improving generative quality over conventional masked LMs. Later, masked diffusion frameworks (Sahoo et al., 2024; Shi et al., 2024) introduced simplified and generalized training objectives that reduce complexity while maintaining performance, and support state-dependent masking strategies. Diffusion Large Language Models (dLLMs) including LLaDA (Nie et al., 2025) and Dream  (Ye et al., 2025) demonstrated that diffusion-based models can scale to billions of parameters, achieving strong context learning and instruction-following capabilities. Techniques such as variance-reduced preference optimization (Zhu et al., 2025) and reinforcement of lateral thought chains (Huang et al., 2025) further enhance generative stability and reasoning ability, while efficient perplexity bounds and ratio matching (Haxholli et al., 2025) improve sampling efficiency. More recently, works such as d1 (Zhao et al., 2025) and wd1 (Tang et al., 2025) have explored the possibility of applying reinforcement learning on top of existing dLLMs. LLaDA 2.0 (Bie et al., 2025) pushed the boundaries of model capacity by scaling diffusion language models to 100B parameters, offering a comprehensive analysis of scaling laws.
5.2 Weighted Methods in Language Model Learning
A complementary line of research focuses on improving large language models through weighted supervised fine-tuning (SFT) and reinforcement learning. Weighted SFT techniques aim to prioritize informative tokens. DFT (Wu et al., 2025) addresses the limited generalization of SFT by dynamically rescaling token-level objectives to stabilize gradient updates. Rho-1 (Lin et al., 2024) challenges the conventional uniform next-token prediction objective by selectively training on useful tokens identified via a reference model. Similarly, Luo et al., 2023 propose a re-weighting strategy for active learning in NER, assigning dynamic smoothed weights to tokens to boost performance. In reinforcement learning, applying weighting to improve learning effectiveness is also a common strategy. Cui et al., 2025 investigate the entropy dynamics of reinforcement learning for reasoning LLMs, and propose entropy-based weighting methods to prevent entropy collapse and encourage exploration. Wang et al., 2025 highlight the critical role of high-entropy minority tokens in RLVR, demonstrating that restricting updates to these tokens significantly enhances reasoning capabilities. Together, these approaches demonstrate that token-level weighting is an important and effective way to improve the optimization of language models.
6 Conclusion
We introduced GIFT, a theoretically grounded importance-aware fine-tuning algorithm for diffusion language models. By leveraging token-level entropy as a measure of importance, GIFT prioritizes high-uncertainty tokens and preserves diffusion process properties. Experiments across varying datasets, different training frameworks, multiple benchmarks, and on both base and instruct models demonstrate the universal advantage of the GIFT method over SFT. These results highlight entropy-based importance weighting as a simple yet powerful strategy for enhancing training effectiveness and stability in diffusion language models.
Limitations
We identify two primary limitations of this work. First, our theoretical derivation relies on certain necessary simplifying assumptions. Specifically, we assume a fixed matrix $Q$ during the derivation; however, in practical entropy-driven masking, $Q$ functions more as a heuristic variance-control scheme. We encourage future research to extend the theoretical framework to account for a dynamic $Q$. Second, the scale of our experimental validation is still limited. This is primarily due to computational constraints. Our computational resources restricted us from using even larger models, more extensive datasets, or a higher number of experimental trials.
Ethical Considerations
Potential Risks.
This paper aims to advance the field of machine learning. The work is solely centered on the methodology itself; how it is applied and for what purpose is entirely up to the users. As with other large-model research, models trained with our algorithm may generate content that is factually incorrect or potentially harmful, and such outputs should be carefully vetted by users.
Usage of Large Language Models.
In this paper, LLMs were used for polishing and improving phrases and sentences, but they did not participate in any aspects of idea conception, experimental design, theoretical proofs, figure creation, or paper structure design.
Appendix A Proof of the Weighted SFT Loss
In order to prove this theorem, we will first prove two lemmas, and then proceed to prove the main theorem. Our proof follows that of Ou et al., 2025, with the key difference that our $Q$ matrix incorporates varying masking rates $\beta$, whereas Ou et al., 2025 assumes all masking rates are identical.
Here, we assume a fixed matrix $Q$ during the derivation; however, in practical entropy-driven masking, $Q$ functions more as a heuristic variance-control scheme. We encourage future research to extend the theoretical framework to account for a dynamic $Q$.
A.1 Proof of Lemmas
Let the original response of the model be $x_{0}$. At a given timestep $t$, let the partially masked response be $x_{t}$, and $[M]$ denotes the mask. $\beta_{x_{t}^{i}}$ denotes the masking rate corresponding to the $i$-th token. The definition of $\bar{f}$ can be found in Equation 6.
Lemma 1.
Let $x_{t}^{i}=[M]$ and ${x^{\prime}}_{t}^{i}\neq[M]$. Then we have
$\frac{p_{t}(x_{t}^{1},\dots,{x^{\prime}}^{i}_{t},\dots,x_{t}^{d})}{p_{t}(x_{t}^{1},\dots,x^{i}_{t},\dots,x_{t}^{d})}$
$\quad=\frac{\exp(-\beta_{x_{t}^{i}}\,\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\,\bar{f}(t))}\;p_{0}({x^{\prime}}_{t}^{i}\mid x_{t}^{\mathrm{unmasked}}).$ (17)
On the other hand, if $x_{t}^{i}\neq[M]$, then
$\frac{p_{t}(x_{t}^{1},\dots,{x^{\prime}}^{i}_{t},\dots,x_{t}^{d})}{p_{t}(x_{t}^{1},\dots,x^{i}_{t},\dots,x_{t}^{d})}=0;$ (18)
Intuitively, this theorem characterizes the exact transition probabilities during the training process. It also establishes a connection between the concrete score function and the clean data distribution via a scalar factor.
Proof.
We first calculate $p_{t|0}(x_{t}|x_{0})$:
Based on the definition of $Q_{t}$, we can easily see that $p_{t|0}(x_{t}|x_{0})=0$ if $x_{t}\neq x_{0}$ and $x_{t}\neq[M]$
Now we divide the time interval $(0,t)$ into $n$ intervals, namely $(0,s_{1}),(s_{1},s_{2}),\cdots,(s_{n-1},t)$. Here we denote $s_{0}=0,s_{n}=t$
then the total probability of $p_{t|0}(x_{t}=x_{0}|x_{0})$ is
$\prod_{i=0}^{n-1}p_{s_{i+1}|s_{i}}(x_{s_{i+1}}=x_{s_{i}}|x_{s_{i}})$
$=\prod_{i=0}^{n-1}\Big(1+Q_{s_{i}}(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})$
$\quad+o(s_{i+1}-s_{i})\Big)$
$=\exp\Big(\sum_{i=0}^{n-1}\ln\Big(1+Q_{s_{i}}(x_{s_{i}},x_{s_{i}})$
$\quad\times(s_{i+1}-s_{i})+o(s_{i+1}-s_{i})\Big)\Big)$
$=\exp\Big(\sum_{i=0}^{n-1}[f(s_{i})Q(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})$
$\quad+o(s_{i+1}-s_{i})]\Big)$
Now we select a series of appropriate $n$ and $s_{1},s_{2},\dots,s_{n-1}$ such that while $n\rightarrow+\infty$ , $\max(s_{i+1}-s_{i})\rightarrow 0$. Now by recalling the definition of integral, we obtain:
$\exp\Big(\sum_{i=0}^{n-1}[f(s_{i})Q(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})$
$\quad+o(s_{i+1}-s_{i})]\Big)\rightarrow\exp(-\beta_{x_{0}}\bar{f}(t))$
Thus, we have
$p_{t|0}(x_{t}|x_{0})=\begin{cases}\exp(-\beta_{x_{0}}\bar{f}(t))\\ \quad\text{if }x_{t}=x_{0}\\ 1-\exp(-\beta_{x_{0}}\bar{f}(t))\\ \quad\text{if }x_{t}=[M]\\ 0\\ \quad\text{otherwise}\end{cases}$
Now, based on the above result, recalling the independence of different dimensions in the diffusion process, we can get a stronger result:
Lemma 2.
Suppose that we have $x_{t}$ in which the tokens $a_{1},a_{2},...,a_{k}$ are masked and $b_{1},b_{2},...,b_{n-k}$ are unmasked; then:
$p_{t}(x_{t})$ $=\prod_{i=1}^{k}(1-\exp(-\beta_{a_{i}}\bar{f}(t)))$
$\quad\times\prod_{j=1}^{n-k}\exp(-\beta_{b_{j}}\bar{f}(t))p_{0}(x_{t}^{\text{unmasked}})$ (19)
Proof.
Exploiting the independence across dimensions in the diffusion process, we may, without loss of generality, set $a_{i}=i$ and $b_{i}=k+i$. Equivalently, this corresponds to masking the first $k$ tokens while leaving the remaining nkn-k tokens unmasked.
Then we obtain
$p_{t}([M]\dots[M]x_{t}^{k+1}\dots x_{t}^{n})$
$=\sum_{x_{0}}p_{t|0}([M]\dots[M]x_{t}^{k+1}\dots x_{t}^{n}|x_{0})p_{0}(x_{0})$
$=\sum_{x_{0}^{1}\dots x_{0}^{n}}\prod_{i=1}^{k}p^{k}_{t|0}([M]|x_{0}^{k})$
$\quad\times\prod_{i=k+1}^{n}p_{t|0}(x_{t}^{k}|x_{0}^{k})p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=\sum_{x_{0}^{1}\dots x_{0}^{n}}\prod_{i=1}^{k}(1-\exp(-\beta_{i}\bar{f}(t)))$
$\quad\times\prod_{i=k+1}^{n}\exp(-\beta_{i}\bar{f}(t))p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=\sum_{x_{0}^{k+1}\dots x_{0}^{n}}\prod_{i=1}^{k}(1-\exp(-\beta_{i}\bar{f}(t)))$
$\quad\times\prod_{i=k+1}^{n}\exp(-\beta_{i}\bar{f}(t))p_{0}(x_{0}^{k+1},\dots x_{0}^{n})$
Recalling the defination, $p_{0}(x_{0}^{k+1},\dots,x_{0}^{n})$ can be expressed as $p_{0}(x_{t}^{\text{unmasked}}),$ the desired result follows.
Now we append $x^{\prime}_{t}$ and $x_{t}$ to the above result and we obtain the result of Lemma 1.
A.2 Proof of the Main Theorem
See 1
Proof.
Now we consider the Denoising Score Entropy, introduced in Lou et al. (2024). DSE loss provides a principled training objective for diffusion models by measuring the discrepancy between the model score function and the true score of the perturbed data distribution.
The DSE loss can be written as:
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}$
$\quad\times\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})\Big(s_{\theta}(x,t)_{x_{t}^{\prime}}$
$\quad-\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}+C\Big)dt$
$C=K\!\left(\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right),\quad K(a)=a\log a-a,$
[…]
In practice, we often estimate $s_{\theta}(x,t)_{x_{t}^{\prime}}$ using
$s_{\theta}(x,t)_{x_{t}^{\prime}}$ $\approx p_{t}(x_{t}^{1}\dots x^{\prime}{}^{i}_{t}\dots x_{t}^{d})$
$\quad\times\frac{1}{p_{t}(x_{t}^{1}\dots x^{i}_{t}\dots x_{t}^{d})}$
We denote $\tilde{Q}$ as the reverse transition matrix corresponding to $Q$. Then, the reverse process satisfies
[…]
Now we use Lemma 1 and calculate the three parts of the loss separately:
Using the fact that
$\sum p_{0}({x^{\prime}}_{t}^{i}|x_{t}^{\text{unmasked}})=1$
we obtain:
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})s_{\theta}(x,t)_{x_{t}^{\prime}}dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}\beta_{i}f(t)$
$\quad\times\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}dt$
Using the properties of $\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}$ we obtain:
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})$
$\quad\times\left(-\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}\right)dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M],j\neq[M]}-f(t)\beta_{i}$
$\quad\times\mathbf{1}[x_{0}^{i}=j]\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}dt$
Recalling that $K(a)=a\log a-a$ with $K(0)=0$, we can safely ignore the terms $\mathbf{1}[x_{0}^{i}=j]$ and $j\neq[M]$.
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}\mid x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})$
$\quad\times K\!\left(\frac{p_{t|0}(x_{t}^{\prime}\mid x_{0})}{p_{t|0}(x_{t}\mid x_{0})}\right)\,dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}\mid x_{0})}\sum_{x_{t}^{i}=[M],j\neq[M]}\beta_{i}f(t)$
$\quad\times K\!\left(\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}})\bar{f}(t)}\mathbf{1}[x_{0}^{i}=j]\right)dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}\mid x_{0})}\sum_{x_{t}^{i}=[M]}\beta_{i}f(t)$
$\quad\times\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}$
$\quad\times\left(\log\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}-1\right)dt$
By adding them all together, especially combining the first and third term, we get
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}-\frac{\beta_{i}f(t)\exp(-\beta_{i}\bar{f}(t))}{1-\exp(-\beta_{i}\bar{f}(t))}$
$\quad\times\log\left(\frac{\exp(-\beta_{i}\bar{f}(t))}{1-\exp(-\beta_{i}\bar{f}(t))}q_{\theta}(x_{0}^{i}|x_{t}^{\text{unmasked}})\right)dt$
Now we apply a change of variables. By defining
$t_{i}=1-\exp\!\big(-\beta_{x_{t}^{i}}\,\bar{f}(t)\big),$
we obtain
$dt_{i}=\beta_{x_{t}^{i}}f(t)\exp\!\big(-\beta_{x_{t}^{i}}\,\bar{f}(t)\big)dt$
and the loss can be rewritten as:
$\sum_{i}\int_{0}^{1}\mathbb{E}_{x\sim p_{t|0}(x_{t}|x_{0})}\Big[-\frac{1}{t_{i}}\mathbf{1}[x_{t}^{i}=[M]]$
$\quad\times\log(q_{\theta}(x_{0}^{i}|x_{t}^{\text{unmasked}}))\Big]dt_{i}+C$
Here, $x_{t_{i}}$ denotes the distribution in which $x^{i}$ is masked with probability $t_{i}$.
Recall that
$t_{i}=1-\exp(-\beta_{x^{i}}\,\bar{f}(t)).$
Suppose we have a reference value $\beta_{\text{ref}}$ and define
$t=1-\exp(-\beta_{\text{ref}}\,\bar{f}(t)).$
Then we can express $t_{i}$ as
$t_{i}$ $=1-\exp(-\beta_{x^{i}}\,\bar{f}(t))$
$=1-\exp(-\beta_{x^{i}}(\frac{1}{\beta_{ref}}\log(1-t)))$
$=1-(1-t)^{\frac{\beta_{x^{i}}}{\beta_{\text{ref}}}}$
In practice, one can sample $t$ and calculate $t_{i}$ using the above formula. ∎
Appendix B Algorithm Pipeline
In Algorithm 1 and Algorithm 2, we provide a detailed description of the implementation of GIFT.
Input: model $\mathcal{M}$, input_ids $\mathcal{I}$, mask_token_id $m$
Output: Masking parameter $\beta$
Function EstimateRate($\mathcal{M},\mathcal{I},M,m$):
// mask out the answer
for $i\leftarrow 1$ to $|\mathcal{I}|$ do
 $\mathcal{I^{\prime}}^{i}\leftarrow(1-M^{i})\cdot\mathcal{I}^{i}+M^{i}\cdot m$;
$z\leftarrow\mathcal{M}(\mathcal{I^{\prime}}).\text{logits}$;
// calculate entropy
for $i\leftarrow 1$ to $|\mathcal{I}|$ do
 $H_{i}\leftarrow-\sum_{j}\text{softmax}(z_{ij})\cdot\text{log\_softmax}(z_{ij})$;
 $\beta_{i}\leftarrow M^{i}\sqrt{H_{i}}$;
return $\beta$;
Algorithm 1 Creating Betas from Entropy
Input: model $\mathcal{M}$, input_ids $\mathcal{I}$, labels $l$, mask_token_id $m$, prompt lengths $L$
Output: Loss $\mathcal{L}$
Function ComputeLoss($\mathcal{M},\mathcal{I},m,L$):
$t\sim\mathcal{U}(0,1)$;
// mask the model’s answer
for $i\leftarrow 1$ to $|\mathcal{I}|$ do
 $M^{i}=\mathbf{1}_{\{\,i>L\,\}}$
$\beta\leftarrow\textnormal{{EstimateRate}}(\mathcal{M},\mathcal{I},M,m)$;
$\beta_{ref}\leftarrow\frac{\sum_{i}\beta_{i}\cdot M_{i}}{\sum_{i}M_{i}}$ ;
// Compute the masking probability for each token
for $i\leftarrow 1$ to $|\mathcal{I}|$ do
 $t_{i}\leftarrow 1-(1-t)^{\frac{\beta_{i}}{\beta_{ref}}}$;
 $M^{\prime i}\sim\text{Bernoulli}(t_{i})$ ;
$\mathcal{I^{\prime}}\leftarrow(1-M^{\prime})\cdot\mathcal{I}+M^{\prime}\cdot m$;
$z\leftarrow\mathcal{M}(\mathcal{I^{\prime}}).\text{logits}$;
$S\leftarrow\sum_{i}\mathbf{1}_{\{M^{\prime i}=1\}}\frac{1}{t_{i}}\cdot$ $\text{CrossEntropy}(z_{i},l_{i})$;
$N\leftarrow\sum_{i}\mathbf{1}_{\{M^{\prime i}=1\}}$;
$\mathcal{L}\leftarrow S/N$;
return $\mathcal{L}$;
Algorithm 2 Weighted Entropy-driven Fine-Tuning
Category Hyperparameter = Value
Optimization optimizer=\ =\ adamw_torch; weight_decay=0.1\ =0.1;
learning_rate = 1e-5; lr_scheduler_type = linear;
adam_beta1 = 0.9; adam_beta2 = 0.999; adam_epsilon = 1e-8;
init_fn=\ =\ mitchell; init_std=0.02\ =0.02; max_grad_norm=1\ =1
LoRA Configuration LoRA=\ =\ True; $r=128$; lora_alpha=256\ =256; lora_dropout=0.05\ =0.05;
target_modules == ["q_proj", "k_proj", "v_proj"]
Training num_train_epochs=20\ =20; per_device_train_batch_size=1\ =1;
gradient_accumulation_steps=4\ =4; precision = bf16;
max_sequence_length = 4096; seed =42\ =42
Table 6: Hyperparameters used in LoRA training.
Category Hyperparameter = Value
Optimization optimizer=\ =\ adamw_torch; weight_decay=0\ =0;
learning_rate = 2e-5; lr_scheduler_type = cosine;
adam_beta1 = 0.9; adam_beta2 = 0.999; adam_epsilon = 1e-8;
init_fn=\ =\ mitchell; init_std=0.02\ =0.02; max_grad_norm=1\ =1
Training num_train_epochs=5\ =5; per_device_train_batch_size=4\ =4;
gradient_accumulation_steps=1\ =1; precision = bf16;
max_sequence_length = 4096; seed =42\ =42
Table 7: Hyperparameters used in full parameter training.
Appendix C Hyperparameters and Detailed Settings
C.1 Training Hyperparameters and Settings
Our training is divided into two settings: LoRA fine-tuning and full parameter fine-tuning. LoRA fine-tuning is conducted on 4 H100 GPUs, while full parameter fine-tuning is conducted on 8 H100 GPUs.
For LoRA fine-tuning, experiments are conducted with a learning rate of $1\times 10^{-5}$ and a total of 20 training epochs. We apply gradient accumulation with $4$ steps and used an effective per-device training batch size of $1$. To reduce memory cost, we employ LoRA with rank $r=128$, scaling factor $\alpha=256$, and a dropout rate of $0.05$. All hyperparameters, including model architecture, optimization, LoRA configuration, and training settings, are listed in Table 6.
For full parameter fine-tuning, we train for 5 epochs with a learning rate of $2\times 10^{-5}$. Other settings are detailed in Table 7.
C.2 Evaluation Hyperparameters and Settings
In the evaluation, we set the block length to 32. For the two smaller datasets, Sudoku and Countdown, the generation length is set to 512; for the two larger datasets, MATH500 and GSM8K, the generation length is set to 256. The number of diffusion steps is always set to half of the generation length, and we use the low-confidence remasking method. Regarding the specific partitioning of validation sets for these four datasets, we follow the settings in wd1 (Tang et al., 2025).
Appendix D Motivation for Using Entropy-based Methods as the Weight Coefficient
In this section, we provide an intuitive perspective to justify the choice of entropy as the coefficient $\beta$, as this section is intended to provide motivation rather than a rigorous proof.
In Alain et al. (2016), the authors proved that importance sampling has minimum variance when the sampling proposal is proportional to the $L_{2}$-norm of the gradient.
Recall that in a model, $p$ is generated by applying softmax to the output logits. More specifically,
$p_{k}=\frac{e^{z_{k}}}{\sum_{j}e^{z_{j}}}$
Here $z$ is the logits generated by the model.
Let $c$ be the label and $p_{c}$ be the probability of this token. Then we have
$\frac{\partial\log p_{c}}{\partial z_{k}}=\left\{\begin{array}[]{ll}-p_{k}&\text{if }k\neq c\\ 1-p_{c}&\text{if }k=c\end{array}\right$.
In the following discussion, we assume that the base model generates the correct token with relatively high probability and other tokens with relatively low probability. Therefore, $L_{2}$-norm of $\frac{\partial\log p_{c}}{\partial\mathbf{z}}$ can be approximated as $1-p_{c}$. Since $\frac{\partial\log p_{c}}{\partial\mathbf{z}}$ constitutes a key component of the loss gradient, theoretical insights suggest that selecting an importance sampling distribution that is positively correlated with this term might help reduce the variance of the loss, thereby enhancing training stability and effectiveness. We now briefly analyze why entropy can serve this purpose.
The entropy function $H(\mathbf{p})$ admits the following lower bound:
$H(\mathbf{p})$ $\geq-p_{c}\log p_{c}+(1-p_{c})$ (20)
$(\log(n-1)-\log(1-p_{c}))$
where $n$ denotes the vocabulary size. Since the term $p_{c}\log p_{c}$ satisfies
$p_{c}\log p_{c}<p_{c}(1-p_{c}),$ (21)
it’s negligible compared to $(1-p_{c})\log(n-1)$. We may safely ignore it for approximation purposes.
Moreover, we empirically observe that the variation in incorrect predictions across different samples is relatively small. Under this assumption, $H(\mathbf{p})$ is positively correlated with $(1-p_{c})$.
On the other hand, using entropy $H(\mathbf{p})$ is superior to directly using $(1-p_{c})$ in that it preserves richer information about the underlying distribution. Due to the prevalence of synonyms and semantically similar tokens, the true probability distribution is unlikely to be unimodal. Instead, it is more reasonable to expect the probability mass to be concentrated on a small subset of tokens. In this respect, entropy $H(\mathbf{p})$ captures such structural information significantly better than $(1-p_{c})$. In pratice, we employ a power function based on $\sqrt{H(\mathbf{p})}$ as the importance weighting function.
Appendix E SFT Role in Subsequent RL
Figure 3: Reward curves of models cold-started with GIFT and SFT during subsequent reinforcement learning training. The curves are smoothed using a time-weighted EMA. As shown, the model initialized with GIFT achieves higher rewards.
Reinforcement learning (RL) post-training has become a crucial stage in the training of large language models  (DeepSeek-AI, 2025). Supervised fine-tuning (SFT) cold-start plays an important role in preparing the model for subsequent RL, as SFT establishes a reliable initial policy which can then be further refined by RL objectives. Therefore, it is essential to verify whether models trained with GIFT can maintain their advantages during subsequent RL training. To this end, we fine-tune models on the s1K dataset using both GIFT and SFT, and then further train them with RL on the Sudoku dataset for comparison.
As shown in Figure 3, models initialized with GIFT achieve higher rewards throughout RL training. Moreover, as reported in Table 8, the GIFT-trained model achieves significantly better evaluation results. These findings demonstrate that models cold-started with GIFT are effective in guiding subsequent RL learning.
Model Base SFT GIFT SFT+RL GIFT+RL
Sudoku 5.5 4.6 7.8 8.9 13.3
Table 8: Evaluation results of models cold-started with GIFT and SFT during subsequent RL training.